OpenStudy (anonymous):

where can i find more info about the Derivation of h(t)=-16t2+vt+h0

OpenStudy (swissgirl):

We are deriving with respect to t \[\frac{d}{dt}(-16t^2+vt+h_0)\] \[\frac{d}{dt}(-16t^2)+\frac{d}{dt}(vt)+\frac{d}{dt}(h_0)\] \[-2*16t+v+0=-32t+v\]

OpenStudy (anonymous):

what is d/dt

OpenStudy (swissgirl):

That means taking the derivative with respect to t

OpenStudy (swissgirl):

It basically just a symbol that means taking the derivative of a function

OpenStudy (swissgirl):

h_0 is a constant and when we take the derivative of a constant the result will be 0 The derivative of a constant is 0 DO you know the rules of derivatives?

OpenStudy (anonymous):

oh, actually i was asking how the formula was made

OpenStudy (anonymous):

like what formulas were used to make this formula

OpenStudy (swissgirl):

hmm dont think I get ur question @ganeshie8 Can you help out?

OpenStudy (anonymous):

how was the formula invented

ganeshie8 (ganeshie8):

familiar with integrals ?

OpenStudy (anonymous):

um, not really...

ganeshie8 (ganeshie8):

that formula gives you the height of a projectile after a given time right ?

OpenStudy (anonymous):

you mean h=-16t2+vt+h?

OpenStudy (swissgirl):

ahhhh I get the question now -.-

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

so, how was the formula made?

ganeshie8 (ganeshie8):

watch this https://www.youtube.com/watch?v=pD5jfwGe7eQ

OpenStudy (anonymous):

ok, thx

ganeshie8 (ganeshie8):

let me know if u have any questions after watching the video

OpenStudy (anonymous):

k